Ann is two and a half inches taller than Tara. There is a four and one fourth inches difference between Tara and Peter. Jamilla is taller than Ann but three fourths inches shoter than Peter. X represents Tara's height - write and simplify in expression involving Jamilla's height.
Answer by ptaylor(2198)

Let X represent Tara's height
Then X+2.5= Ann's height (we are told this)
X+4.25=Peter's height
(How do we know this? Jamilla is taller than Ann and therefore taller than Tara but Jamilla is 0.75 inches shorter than Peter. So Peter has to be taller than Tara----by 4.25 inches)
So (X+4.25)-0.75=Jamilla's height. Simplifying, we get:
(X+3.5)=Jamilla's height

Hope this helps----ptaylor